A conductor possesses a positive charge of 3.2x 10^-19 C. How many electron does it have in excess or deficit (e = 1.60 x 10^-19C)
a.2
b.3
c.5
d.6
1
Expert's answer
2019-04-01T10:46:03-0400

We can find the number of the electrons in the conductor from the formula:


"q = Ne,"

here, "q" is the charge of the conductor, "N" is the number of electrons in the conductor, "e" is the charge of the electron.

Then, we get:


"N = \dfrac{q}{e} = \dfrac{3.2 \cdot 10^{-19}C}{1.6 \cdot 10^{-19} C} = 2."

Since the charge of the conductor is positive, the conductor has a deficit of 2 electrons.

Answer:

The conductor has a deficit of 2 electrons.